# Divide-and-Conquer Meets Consensus: Unleashing the Power of Functions in Code Generation

### v1

@@ A.6 Supplementary Results
We provide token usage results in Table 11 for FunCoder and baseline methods on the HumanEval dataset with the GPT-3.5 model. We report the average token usage per problem. The token usage is computed through the sum of prompt tokens and completion tokens returned by OpenAI API chat completion call [^4] https://platform.openai.com/docs/guides/text-generation/managing-tokens. For LDB, we report their token usage in the original paper (Zhong et al., 2024).
@@
Table 11: Token usage for different settings of FunCoder and baseline methods. The LDB results are reported in the original paper. The main setting for LDB and FunCoder is bolded.
Method Setting Pass@1 Avg. Tokens
Standard One-time 68.3 886.7
CodeT One-time + Self-Test@11 81.1 (+12.8) 4479.1
Reflexion maxiter=2 69.5 (+1.2) 1416.1
LDB (reported) line-level 80.5 (+12.2) 24K
block-level 82.9 (+14.6) 23K
function-level 79.9 (+11.6) 27K
FunCoder One-pass 72.6 (+4.3) 1233.7
Two-pass 78.7 (+10.4) 3343.2
Two-pass + Consensus@5 83.5 (+15.2) 4040.8
Two-pass + Consensus@11 85.4 (+17.1) 5402.0
FunCoder (self-test ablation) Two-pass + Self-Test@11 80.5 (+12.2) 5408.3
Full Results for MATH Levels
@@
The MATH dataset divides the problems into five levels of difficulty. The difficulty distribution of our test set can be found in Table 8. We report the average accuracy of FunCoder and other methods at each level in Table 12. The results of Cumulative Reasoning are obtained from the original paper (Zhang et al., 2024). Experiment results demonstrate that our method consistently enhances the model’s reasoning ability across all levels of MATH.
@@
GPT-3.5 Standard 61.5 51.1 43.5 25.9 17.7 34.6
CoT 76.9 48.9 50.9 33.6 21.8 40.0
PoT 61.5 51.1 56.5 33.6 24.1 41.0
Self-Refine 84.6 61.1 65.7 32.8 31.3 48.6
FunCoder 84.6 65.9 68.5 43.1 37.4 54.0

### v2

@@ A.9 Supplementary Results
We provide token usage results in Table 9 for FunCoder and baseline methods on the HumanEval dataset with the GPT-3.5 model, whereas usage results on other datasets are provided in Table 10. We report the average token usage per problem. The token usage is computed through the sum of prompt tokens and completion tokens returned by OpenAI API chat completion call [^4] https://platform.openai.com/docs/guides/text-generation/managing-tokens. For LDB, we report their token usage in the original paper (Zhong et al., 2024).
@@
Table 9: Token usage for different settings of FunCoder and baseline methods on HumanEval, all evaluated on GPT-3.5-turbo. The LDB results are reported in the original paper. The main setting for LDB and FunCoder is bolded.
Method Setting Pass@1 Tokens
Min. Max. Avg. Med.
Standard One-time 68.3 648 1477 886.7 861
CodeT One-time + Self-Test@11 81.1 (+12.8) 2298 9645 4479.1 4166
Reflexion maxiter=2 69.5 (+1.2) 416 4906 1416.1 754
LDB (reported) line-level 80.5 (+12.2) - - 24K -
block-level 82.9 (+14.6) - - 23K -
function-level 79.9 (+11.6) - - 27K -
FunCoder One-pass 72.6 (+4.3) 826 3489 1233.7 1132
Two-pass 78.7 (+10.4) 2197 8406 3343.2 3078
Two-pass + Consensus@5 83.5 (+15.2) 2455 9432 4040.9 3800
Two-pass + Consensus@11 85.4 (+17.1) 3015 13850 5402.0 5166
FunCoder Two-pass + Self-Test@11 80.5 (+12.2) 2967 13758 5408.3 5184
(ablation) Two-pass + Clustering@11 75.0 (+6.7) 3044 9958 5070.7 4888
Table 10: Token usage of FunCoder and baseline methods on other datasets, i.e. MBPP, xCodeEval and MATH. Results are evaluated on GPT-3.5-turbo.
Dataset Method Pass@1 Tokens
Min. Max. Avg. Med.
MBPP Standard 72.0 577 2102 744.9 717.0
CodeT 76.0 (+4.0) 2232 8172 2945.3 2866.0
Reflexion 72.5 (+0.5) 391 3379 1205 569.5
FunCoder 78.5 (+6.5) 3462 13229 5049.9 4644.0
xCodeEval Standard 20.2 1051 3343 1599.5 1530.0
CodeT 23.2 (+3.0) 2264 9245 3937.4 3733.0
Reflexion 20.6 (+0.4) 2977 1003222 401767.3 328591.5
FunCoder 31.4 (+11.2) 4883 53225 10559.7 8927.0
MATH PoT 41.0 551 5867 953.0 835.0
FunCoder 54.0 (+13.0) 2622 30139 7075.5 5666.5
Full Results for Code Generation
We provide results for all conducted experiments on code generation benchmarks in Table 11. Our method consistently improves the baseline on community models by averaging $11\%$ on MBPP and $150\%$ on xCodeEval. It is worth noting that small models have a tendency to have low pass rates on competition problems, leading to a relatively higher randomness, therefore we run 3 experiments and report the median result.
Table 11: Results for Code Generation. We report Pass@1 as evaluate metric. Results from the original paper are underlined, and the best results are bold.
Model Method HumanEval MBPP xCodeEval
GPT-3.5 Standard 68.3 - 72.0 - 44.4 15.2 4.6 0.0 20.2
CodeT 81.1 +12.8 76.0 +4.0 50.6 16.1 8.0 0.0 23.2
Reflexion 69.5 +1.2 72.5 +0.5 44.4 17.0 5.7 0.0 20.6
LDB 82.9 +14.6 76.0 +4.0 - - - - -
FunCoder 85.4 +17.1 78.5 +6.5 62.4 29.5 11.6 0.0 31.4
GPT-4 Standard 82.9 - 73.5 - 68.5 39.3 19.5 1.7 37.4
Parsel 85.0 +2.1 - - - - - - -
CodeT 90.9 +8.0 77.0 +3.5 76.4 51.8 21.8 3.4 44.0
Reflexion 91.0 +8.1 77.1 +3.6 71.3 41.1 19.5 2.5 38.6
MetaGPT 85.9 +3.0 - - - - - - -
FunCoder 94.5 +11.6 79.5 +6.0 83.1 58.0 26.4 3.4 48.6
GPT-4o mini Standard 87.3 - 76.0 - 65.7 44.6 9.2 0.0 35.4
CodeT 90.9 +3.6 75.5 -0.5 71.9 49.1 16.1 0.0 39.6
FunCoder 91.5 +4.2 77.5 +1.5 72.5 52.3 11.5 0.0 39.8
Llama38b Standard 61.6 - 60.5 - 9.0 1.8 0.0 0.0 3.6
CodeT 68.9 +7.3 61.5 +1.0 12.4 0.0 0.0 0.0 4.4
FunCoder 79.7 +18.1 62.5 +2.0 22.0 0.9 0.0 0.0 8.0
Codestral22b Standard 79.3 - 68.5 - 27.5 4.5 2.3 0.0 11.4
CodeT 86.0 +7.3 74.0 +5.5 34.8 7.1 3.4 0.0 14.8
FunCoder 89.0 +9.7 74.5 +6.0 49.4 15.2 3.4 0.0 22.0
StableCode3b Standard 61.0 - 51.5 - 7.3 0.9 0.0 0.0 2.8
CodeT 75.0 +14.0 57.5 +6.0 11.2 1.8 0.0 0.0 4.6
FunCoder 81.0 +20.0 63.5 +12.0 13.5 4.5 1.1 0.0 6.2
CodeLlama34b Standard 43.9 - 53.5 - 2.3 0.0 0.0 0.0 0.8
CodeT 55.5 +11.6 56.5 +3.0 10.1 0.0 0.0 0.0 3.6
FunCoder 66.5 +22.6 58.5 +5.0 10.2 0.0 0.0 0.0 3.6
StarCoder215b Standard 59.8 - 64.5 - 18.0 0.9 2.3 0.0 7.2
CodeT 70.7 +10.9 66.0 +1.5 13.5 0.9 0.0 0.0 5.0
FunCoder 78.7 +18.9 70.0 +5.5 29.2 4.5 0.0 0.0 11.6
Full Results for MATH
@@
The MATH dataset divides the problems into five levels of difficulty. The difficulty distribution of our test set can be found in Table 7. We report the average accuracy of FunCoder and other methods for each math subject in Table 12 and results for each level in Table 13. The results of Cumulative Reasoning are obtained from the original paper (Zhang et al., 2024). Experiment results demonstrate that our method consistently enhances the model’s reasoning ability across all levels of MATH.
Model Method Prealg. Alg. NT Prob. Geo. InterAlg. Precalc. Overall
GPT-3.5 Standard 62.2 37.4 20.0 29.8 31.0 24.4 21.8 34.6
CoT 59.8 51.1 28.9 29.8 28.6 26.7 30.9 40.0
PoT 68.3 50.4 33.3 48.9 21.4 18.2 29.1 41.0
Self-Refine 74.4 49.6 48.9 57.4 28.6 35.6 36.4 48.6
FunCoder 76.8 61.2 55.6 59.6 34.1 36.0 41.8 54.0
GPT-4 Standard 81.7 82.7 71.1 72.3 59.5 46.7 47.3 68.2
CoT 84.1 87.1 62.2 68.1 45.2 48.9 54.5 68.6
PoT 79.3 80.6 75.6 72.3 50.0 47.8 58.2 68.2
Self-Refine 82.9 82.0 77.8 76.6 54.8 55.6 63.6 72.2
CR 86.6 86.3 88.7 71.1 53.7 51.5 51.8 72.2
FunCoder 89.0 92.8 82.2 83.0 59.5 63.3 56.4 78.2
GPT-4o mini Standard 79.3 83.5 75.6 87.2 47.6 57.8 56.4 71.8
CoT 90.2 95.7 82.2 68.1 50.0 61.1 61.8 77.2
PoT 80.5 84.2 77.8 72.3 50.0 60.0 50.9 71.0
Self-Refine 79.3 83.5 75.6 87.2 47.6 57.8 56.4 71.8
FunCoder 81.7 83.5 80.0 80.9 59.5 60.0 54.5 73.2
Llama38b CoT 56.1 47.5 31.1 34.0 40.5 14.4 38.2 38.6
PoT 67.1 32.4 24.4 34.0 16.7 21.1 18.2 32.6
FunCoder 67.9 45.7 51.1 53.2 19.0 37.8 30.9 45.0
Codestral22b PoT 70.7 56.1 46.7 44.7 21.4 26.7 30.9 45.6
FunCoder 81.7 61.9 46.7 55.3 28.6 45.6 38.2 54.8
StableCode3b PoT 20.7 14.4 17.8 25.5 4.8 8.9 9.1 14.4
FunCoder 46.3 30.2 20.0 29.8 4.8 20.0 18.2 26.6
CodeLlama34b PoT 35.5 26.1 15.0 16.7 0.0 5.5 33.3 15.2
FunCoder 44.8 46.1 37.8 34.1 13.6 24.6 37.5 24.4
StarCoder215b PoT 46.3 29.5 28.9 25.5 21.4 27.8 23.6 30.2
FunCoder 72.0 39.6 40.9 46.8 23.8 28.1 27.3 40.8
@@
GPT-4o mini Standard 87.2 82.2 80.6 62.9 61.9 71.8
CoT 97.4 90.0 87.0 71.6 61.2 77.2
PoT 89.7 81.1 76.9 63.8 61.2 71.0
Self-Refine 87.2 82.2 80.6 62.9 61.9 71.8
FunCoder 94.9 82.2 81.5 62.9 63.9 73.2